OpenStudy (anonymous):

Algebra I 9B Help!! 9/3x - 1 subtracted by 5x/2x + 3

OpenStudy (anonymous):



OpenStudy (anonymous):

I know its sloppy but its hard on the computer... Hopfully you can read it and show me how to workit out.

OpenStudy (anonymous):

\[\frac{9}{3x-1}-\frac{5}{2x+3}\]?

OpenStudy (anonymous):

Yes thats correct @satellite73

OpenStudy (anonymous):

oops another typo \[\frac{9(2x+3)-5x(3x-1)}{(3x-1)(2x+3)}\]

OpenStudy (anonymous):

you can clean up the numerator by using the distributive property and then combining like terms

OpenStudy (anonymous):

@satellite73 Can you help me with this step,, i am not any good at this stuff.